The general Equation of a conic section is ax^2+2hxy+by^2+2gx+2fy+c=0;

Then how did we get the discriminant as abc+2fgh-af^2-bg^2-ch^2 ??

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Moni

The equation ax2 + 2hxy + by2 + 2gx + 2fy + c = 0. Represents a second degree equation where a, h, b doesn’t variables simultaneously.
 
Let a ≠ 0.
 
Now, the above equation becomes
 
        a2 x2 + 2ax (hy + g) = aby2 – 2afy – ac
 
on completing the square on the left side, we get,
 
        a2 x2 + 2ax (hy + g) = y2 (h2 – ab) + 2y (gh – af) + g2 – ac.
 
i.e.    (ax + hy + g) = + √y2(h2–ab)+2y(gh–af) +(g2–ac)
 
We cannot obtain x in terms of y, involving only terms of the first degree, unless the quantity under the radical sign be a perfect square. The condition for this is,
 
(gh – af)2 = (h2 – ab) (g2 – ac)
 
i.e. g2h2 – 2afgh + a2f2 = g2h2 – abg2 – abg2 – ach2 + a2bc
 
        cancelling and diving by a, we have the required condition
 
                abc + 2fgh – af2 – af2 – bg2 – ch2 = 0

Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here
and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.


 All the best.


Regards,
Askiitians Experts
Badiuddin